# Multimodal Laryngoscopic Video Analysis for Assisted Diagnosis of Vocal Fold Paralysis

### v2

@@ 4 Glottis Image Segmentation
However, the U-Net model tends to produce false positives, especially in cases where the glottal area is not visible (see Table 5). These incorrect segmentation can complicate subsequent vocal fold analysis modules. To address this issue and enhance the precision of the segmentation results, we further refine the U-Net outputs using a diffusion model. This additional step helps to correct the errors and produce a more accurate and reliable glottis mask. The combination of these two methods ensures high-quality segmentation, which is crucial for subsequent analysis and metric computation.
@@ 4.1 U-Net-based Method
There are several advantages to using a U-Net model with this structure. First, the architecture captures both low-level and high-level features through its deep structure and skip connections, which concatenate features from the contracting path to the corresponding layers in the expansive path. This facilitates precise segmentation by preserving spatial information. Second, the use of ConvBlocks with batch normalization stabilizes and accelerates the training process, while ReLU activation functions introduce non-linearity, enabling the model to learn complex patterns. Additionally, the upsampling layers allow for finer segmentation outputs by gradually reconstructing the image details.
@@ 4.2 Diffusion Model-based Refinement
The U-Net model excels at extracting image masks. However, the generated masks may tend to yield false alarms (see Table 5). To further enhance the quality, we explore the integration of the diffusion model. Diffusion models are generative models, which proves to be very powerful tools for generating high-quality images [44, 45]. These models are designed to generate images by iteratively denoising random noise to match a target distribution, allowing them to produce diverse outputs while still maintaining a coherent structure or theme. By incorporating the diffusion model as a post refinement step, we expect the glottis mask extracted by U-Net to be further optimized with the diversity introduced.
@@
Consistent with the conventional implementation of the Diffusion Probability Model (DPM) [44], a U-Net model is employed for training. As Fig. 5 is shown, following the idea of MedSegDiff [46], we incorporate the original glottis images as priors for the step estimation function and employ dynamic conditional encoding to fuse the encoding outcomes from both the raw image and the segmentation mask at each step. Hence, for each step, the estimation function $\epsilon_{\theta}$ is written as:
@@ 5 Multimodal Vocal Fold Paralysis Analysis
To detect VFP, we adopt a two-stage process. First, we build a binary classification model to check if it belongs to VFP. Then, with our proposed metrics, left and right VFP can be identified by comparing the left and right vocal fold movement via the plots.
5.1 Vocal Fold Dynamic Extraction with Quadratic Fitting
To diagnose the laryngeal paralysis in a more detailed way, instead of AGA or GAW that measures the both left and right vocal folds, we evaluate the angle deviation for each vocal fold. As a result, we can provide the contrastive diagnosis of unilateral laryngeal paralysis. The procedure for extracting glottal angles from a segmented glottis involves a systematic series of steps, detailed in Algorithm 1, with visual representations of the intermediate results provided in Fig. 6.
Initially, the algorithm acquires the coordinates of the top, bottom, left, and right vertices from the glottis mask, computing their center point denoted as $U(x_{1},y_{1})$, $D(x_{2},y_{2})$, $L(x_{3},y_{3})$, $R(x_{4},y_{4})$, and $C(x_{c},y_{c})$ respectively (see Fig. 6(a)). Subsequently, a line connecting points $C$ and $D$ is established, and a linear function $f(x)$ passing through $C$ and $D$ is computed.
Equidistant points $C_{1},C_{2},...,C_{N-1}$ are then positioned along the line segment intercepted by the glottal mask. For each $C_{k}$, orthogonal functions $f^{\prime}_{k}(x)$ to $f(x)$ are calculated, with their intersection points $L_{k}$ and $R_{k}$ with the glottis mask determined for $k\in[1,N-1]$ (see Fig. 6(b)).
To ensure uniformity, the coordinate system is rotated by an angle $\gamma$, aligning all intersection points along the y-axis after rotation, denoted as $L_{\gamma,k}$ and $R_{\gamma,k}$. Leveraging these points, a quadratic curve $q_{\gamma}(x)$ is approximated in the rotated coordinate system (see Fig. 6(c). The lowest point $D_{q}$ of $q_{\gamma}(x)$ is identified and mapped back to the original coordinate system (see Fig. 6(d)).
@@ 5.1 Vocal Fold Dynamic Extraction with Quadratic Fitting
5.2 Multimodal Vocal Fold Paralysis Detection
We developed a model that integrates both audio and video modalities for the enhanced VFP detection. As illustrated in Fig. 7, the model utilizes the audio spectrograms and the Vocal Fold Dynamics (VFDyn) as inputs. The audio spectrogram is encoded using EfficientNet-b0 [47], a compact and efficient model renowned for its performance in image classification tasks. Given that multiple video highlights are extracted from a single laryngoscope video, with corresponding VFDyn for each, multiple time series are generated. These time series are treated as multi-channel inputs to a ConvLSTM [48] model, comprising a convolutional layer followed by an LSTM layer. The ConvLSTM model is adept at handling multi-channel time series; the convolutional layer captures features across all channels, while the LSTM layer processes temporal information effectively.
To further differentiate between left or right VFP, we analyze the variance of the LVFDyn and the RVFDyn time series. Intuitively, the paralyzed side exhibits less activity during phonation, leading to a smoother time sequence and consequently a lower variance. Therefore, by simply comparing the variance of left and right, we should be able to distinguish between left or right VFP.
6 Experimental Results
@@ 6.2 SYSU Dataset
This dataset was collected in real-world clinical settings by the Sun Yat-sen Memorial Hospital of Sun Yat-sen University (SYSU). This study was approved by the Institutional Review Board of Sun Yat-sen Memorial Hospital of Sun Yat-sen University under no. SYSEC-KY-KS-2022-040 and the Institutional Review Board of Duke Kunshan University under no. 2022ML067. The dataset is deliberately divided into two parts, SYSU-A and SYSU-B.
The SYSU-A dataset consists of 520 clinical video samples, including 106 normal cases and 414 paralysis cases (257 left VFP and 157 right VFP), and is used to evaluate the performance of the proposed MLVAS in UVFP detection. To ensure fairness, we deliberately select equal number of samples for each category for testing. Specifically, we randomly choose 75 samples, including 25 normal, 25 left VFP and 25 right VFP, for testing, and leave other samples for training. The multi-modal VFP detection model is trained and tested on the SYSU-A dataset.
The SYSU-B dataset contains clinical video samples from cases other than VFP, e.g. vocal fold polyps and vocal fold nodules. We randomly split the SYSU-B dataset into training and testing subsets, assigning 80% of the samples for training and 20% for testing. The audio KWS model is trained and tested on the SYSU-B dataset, and its performance is then evaluated on the SYSU-A dataset. It should be noted that there is no overlap between the SYSU-A and SYSU-B datasets, ensuring the model can be applied to the full SYSU-A dataset for evaluation.
All videos were recorded using laryngeal videostroboscopes, with each clip containing at least one strobing video segment. The dataset reflects real-world variability, as the recordings were captured using different types of endoscopes across various locations, leading to domain shifts and posing challenges for algorithms to address.
@@
Dataset # of samples Accuracy(%) Precision(%) Recall(%) F-score(%)
@@ 6.3 Keyword Spotting Model
The detailed architecture of the KWS model is outlined in Table 1. The audio input is represented by a Mel-spectrogram with 80 Mel filters, utilizing 1024 FFT points and a hop length of 512. A sliding window of 400 samples with a hop length of 64 samples is employed to extract frame-level features. During training, audio clips are randomly sampled from the SYSU-A dataset. For each selected audio clip, random chunks of 40 frames are extracted to form batches of training data. The model is trained over 100 epochs using the Adam optimizer with a learning rate of 0.005. During inference, a sliding window of 40 frames with a unit step size is used to process each audio clip in the test dataset.
Since the KWS model outputs a posterior score rather than a direct decision value, a threshold is necessary to determine the final result. Consequently, the threshold is selected based on the highest F-score, a balanced metric that considers both precision and recall, providing a comprehensive measure of classification performance. In MLVAS, we pick a threshold of 0.38 as the classification boundary. If the posterior score output by the KWS model is higher than this value, the input chunk is regarded as ”e”. Otherwise, it is treated as other voices. Fig. 8(a) and Fig. 8(b) display the confusion matrix computed on the SYSU-A and SYSU-B dataset respectively. The accuracy, precision, recall and F-score are listed in Table. 4. As is shown in Table. 4, the performance of our audio KWS model are all above 90% on both the SYSU-A and SYSU-B.
@@
Table 6: Different system settings. The proposed system contains all the modules and uses all the modalities. System 1 is the naive system, using the images segmented by the U-Net baseline without quadratic fitting or diffusion refinement. System 2 adds the quadratic fitting module to the naive system. System 3 and System 4 are two systems with both proposed modules but using only one modality.
System Quadratic Fitting Diffusion-refinement AGA Modality Audio Modality
Proposed System
System 1 - -
System 2 -
System 3 -
System 4 -
Table 7: The average performance and its standard deviation (reported within the parenthesis) across the experiments using 4 different random seeds. All the experiments are conducted on the SYSU-A dataset.
Systems Accuracy (%) Precision (%) Recall (%) F-score (%) Sensitivity (%) Specificity (%)
System 1 73.00 ($\pm$2.45) 72.33 ($\pm$4.63) 76.00 ($\pm$8.00) 73.83 ($\pm$2.48) 76.00 ($\pm$8.00) 70.00 ($\pm$10.04)
System 2 77.67 ($\pm$1.51) 78.00 ($\pm$2.76) 77.33 ($\pm$4.13) 77.67 ($\pm$1.63) 77.33 ($\pm$4.13) 78.00 ($\pm$4.20)
System 3 73.67 ($\pm$2.34) 73.00 ($\pm$3.69) 75.33 ($\pm$1.63) 74.33 ($\pm$1.51) 75.33 ($\pm$1.63) 72.00 ($\pm$5.66)
System 4 66.33 ($\pm$4.63) 65.33 ($\pm$3.56) 69.33 ($\pm$12.56) 67.00 ($\pm$6.90) 69.33 ($\pm$12.56) 63.33 ($\pm$6.41)
Proposed System 81.33 ($\pm$3.27) 81.00 ($\pm$2.28) 82.00 ($\pm$7.04) 81.50 ($\pm$3.73) 82.00 ($\pm$7.04) 80.67 ($\pm$3.01)
@@ 6.5 Vocal Fold Paralysis Detection
We performed ablation studies using various system configurations to evaluate the effectiveness of our proposed methods, as shown in Table 6. The experiments primarily focus on two aspects: first, the effectiveness of the modules for improving metric extraction, such as quadratic fitting and diffusion refinement, and second, the impact of multimodality on VFP detection performance. To validate the statistical significance of the performance improvements observed between different system configurations, a series of t-tests is conducted across different experimental trials with different random seeds. These pairwise t-tests are performed among various system configurations (shown in Table 8), providing a p-value as a measure of statistical significance for the observed differences in performance metrics.

### v3

@@ 4 Visual Feature Extraction with Enhanced Glottis Image Segmentation
However, the U-Net model tends to produce false positives [43, 44, 45], especially in cases where the glottal area is not visible (see the statistics shown in Table 5), resulting in segmentation outputs even in the absence of glottal regions. This can significantly impair the accuracy of subsequent analyses, and complicate subsequent vocal fold analysis modules. To address this issue and enhance the precision of the segmentation results, we further refine the U-Net outputs using a diffusion model. This additional step helps to correct the errors and produce a more accurate and reliable glottis mask. The combination of these two methods ensures high-quality segmentation, which is crucial for subsequent analysis and metric computation.
@@ 4.2 Diffusion Model-based Refinement
The U-Net model excels at extracting image masks. However, the generated masks may tend to yield false alarms when no glottis area is shown (see Table 5). To further enhance the quality, we explore the integration of the diffusion model. Diffusion models are generative models, which proves to be very powerful tools for generating high-quality images [46, 47]. These models are designed to generate images by iteratively denoising random noise to match a target distribution, allowing them to produce diverse outputs while still maintaining a coherent structure or theme. By incorporating the diffusion model as a post refinement step, we expect the glottis mask extracted by U-Net to be further optimized with the diversity introduced.
@@
Consistent with the conventional implementation of the Diffusion Probability Model (DPM) [46], a U-Net model is employed for training. As Figure 4 is shown, following the idea of MedSegDiff [48], we incorporate the original glottis images as priors for the step estimation function and employ dynamic conditional encoding to fuse the encoding outcomes from both the raw image and the segmentation mask at each step. Hence, for each step, the estimation function $\epsilon_{\theta}$ is written as:
@@
4.3 Vocal Fold Dynamics (VFDyn) Extraction with Quadratic Fitting
To diagnose the laryngeal paralysis in a more detailed way, instead of AGA or GAW that measures the both left and right vocal folds, we evaluate the angle deviation for each vocal fold. As a result, we can provide a more comprehensive diagnosis of UVFP. The procedure for extracting glottal angles from a segmented glottis involves a systematic series of steps, detailed in Algorithm 1, with visual representations of the intermediate results provided in Figure 5.
Initially, the algorithm acquires the coordinates of the top, bottom, left, and right vertices from the glottis mask, computing their center point denoted as $U(x_{1},y_{1})$, $D(x_{2},y_{2})$, $L(x_{3},y_{3})$, $R(x_{4},y_{4})$, and $C(x_{c},y_{c})$ respectively (see Figure 5(a)). Subsequently, a line connecting points $C$ and $D$ is established, and a linear function $f(x)$ passing through $C$ and $D$ is computed.
Equidistant points $C_{1},C_{2},...,C_{N-1}$ are then positioned along the line segment intercepted by the glottal mask. For each $C_{k}$, orthogonal functions $f^{\prime}_{k}(x)$ to $f(x)$ are calculated, with their intersection points $L_{k}$ and $R_{k}$ with the glottis mask determined for $k\in[1,N-1]$ (see Figure 5(b)).
To ensure uniformity, the coordinate system is rotated by an angle $\gamma$, aligning all intersection points along the y-axis after rotation, denoted as $L_{\gamma,k}$ and $R_{\gamma,k}$. Leveraging these points, a quadratic curve $q_{\gamma}(x)$ is approximated in the rotated coordinate system (see Figure 5(c). The lowest point $D_{q}$ of $q_{\gamma}(x)$ is identified and mapped back to the original coordinate system (see Figure 5(d)).
@@ 5 Multimodal Vocal Fold Paralysis Analysis
To detect VFP, we adopt a two-stage process. First, we build a binary classification model to check if it belongs to VFP. Then, with our proposed metrics, left- and right-sided VFP can be identified by comparing the left and right vocal fold movement, which is measured by the extracted LVFDyn and RVFDyn features.
@@
We developed a model that integrates both audio and visual modalities for the enhanced VFP detection. As illustrated in Figure 6, the model utilizes the audio Mel spectrograms and the VFDyn as inputs. The audio spectrogram is encoded by the Dasheng audio encoder. Given that multiple key video segments are extracted from a single laryngoscope video, each associated with corresponding VFDyn features, multiple time series are generated for one video. As described in Section 4.3 that multiple equidistant points are selected along the vocal folds, these time series are treated as multi-channel inputs to a lightweight ConvLSTM [49] model, which consists of a convolutional layer followed by an LSTM layer. The ConvLSTM model is well-suited for handling multi-channel time series data: the convolutional layer captures spatial features across all channels, while the LSTM layer effectively processes temporal information. The resulting audio embedding $e_{1}$ and video embedding $e_{2}$ are then concatenated to form a combined embedding $e$. Finally, a classification layer is added to the model, which outputs the logits for VFP prediction.
To further differentiate between left or right VFP, we compute and compare the variance of the LVFDyn and RVFDyn features. Intuitively, the paralyzed side exhibits less activity during phonation, leading to a smoother time sequence and consequently a lower variance. Therefore, by simply comparing the variance of left and right, we are able to distinguish between left or right VFP.
6 Experimental Settings
@@ 6.2 SYSU Dataset
This dataset was collected in real-world clinical settings by the Sun Yat-sen Memorial Hospital of Sun Yat-sen University (SYSU). This study was approved by the Institutional Review Board of Sun Yat-sen Memorial Hospital of Sun Yat-sen University under no. SYSEC-KY-KS-2022-040 and the Institutional Review Board of Duke Kunshan University under no. 2022ML067, 2023LM161 and 2025LM041. The dataset is deliberately divided into two parts, SYSU-A and SYSU-B.
The SYSU-A dataset consists of 520 clinical video samples, including 106 normal cases and 414 paralysis cases (257 left VFP and 157 right VFP), and is used to evaluate the performance of the proposed MLVAS in UVFP detection. To ensure fairness, in the following experiment, we conduct a stratified 10-fold cross validation on the original dataset, ensuring the ratio between positive and negative samples to be the same for the training and testing dataset.
The SYSU-B dataset contains clinical video samples from cases other than VFP, e.g. vocal fold polyps and vocal fold nodules. We randomly split the SYSU-B dataset into training and testing subsets, assigning 80% of the samples for training and 20% for testing. The audio KWS model is trained and validated on the SYSU-B dataset, and its performance is then further evaluated on the SYSU-A dataset. It should be noted that there is no video overlap between the SYSU-A and SYSU-B datasets, ensuring the model can be applied to the full SYSU-A dataset for evaluation.
All videos are recorded using laryngeal videostroboscopes, with each clip containing at least one strobing video segment. The dataset reflects real-world variability, as the recordings were captured using different types of endoscopes across various locations, leading to domain shifts and posing challenges for algorithms to address.
(a)     SYSU-B-test set
(b)     SYSU-A set
@@
Dataset # of samples Accuracy Precision Recall F-score
@@ 6.3 Keyword Spotting Model
The detailed architecture of the KWS model is outlined in Table 1. The audio input is represented by a Mel-spectrogram with 80 Mel filters, utilizing 1024 FFT points and a hop length of 512. A sliding window of 400 samples with a hop length of 64 samples is employed to extract frame-level features. During training, audio clips are randomly sampled from the SYSU-B dataset. For each selected audio clip, random chunks of 40 frames are extracted to form batches of training data. The model is trained over 100 epochs using the Adam optimizer with a learning rate of 0.005. During inference, a sliding window of 40 frames with a unit step size is used to process each audio clip in the test dataset.
Since the KWS model outputs a posterior score rather than a direct decision value, a threshold is necessary to determine the final result. Consequently, the threshold is selected based on the highest F-score, a balanced metric that considers both precision and recall, providing a comprehensive measure of classification performance. In MLVAS, we pick a threshold of 0.38 as the classification boundary. If the posterior score output by the KWS model is higher than this value, the input chunk is regarded as ”e”. Otherwise, it is treated as other voices. Figure 7(a) and Figure 7(b) display the confusion matrix computed on the SYSU-B-test and SYSU-A dataset respectively. The accuracy, precision, recall and F-score are listed in Table 4. As is shown in Table 4, the performance of our audio KWS model are all above 90% on both the SYSU-A and SYSU-B.
@@ 6.5 Vocal Fold Paralysis Detection
The multimodal VFP detection model consider both audio and video modalities. The input for audio modality is 64-dimensional log-Mel spectrogram, extracted every 10 ms with a window size of 32 ms for clips of length 10 s. RVFDyn and LVFDyn are used for the video modalities. As for the models, we use Dasheng-Base that has a model size of 86 M parameters as the backbone pre-trained model for audio modeling. ConvLSTM (described in Section 5) is used for video modeling.
During training, we adopt a full fine-tuning strategy with the learning rate set to $5\times 10^{-5}$. The models are trained for 800 steps using a batch size of 12, comprising 8 positive samples and 4 negative samples. As each laryngeal video typically contains 2–3 key video segments, one segment is randomly selected from each video for training, ensuring that each batch includes video segments from 12 unique laryngeal videos. During inference, class logits are computed for each key segment within a laryngeal video, and the final logits for the video are obtained by averaging the logits across all its key segments.
Table 6: The average classification performance (%) achieved using different audio-based methods, along with the standard deviation (reported in parentheses), is evaluated across experiments using 10-fold cross-validation on the SYSU-A dataset.
Model Feature ROC-AUC Accuracy Precision Recall F-score Sensitivity Specificity
Mean of
Sensi. & Speci.
RF [20] MFCC 78.59($\pm$7.84) 56.23($\pm$6.42) 53.68($\pm$4.17) 99.27($\pm$1.56) 69.56($\pm$3.10) 99.27($\pm$1.56) 13.18($\pm$13.87) 56.23($\pm$6.42)
SVM MFCC 74.81($\pm$7.56) 68.41($\pm$6.75) 70.88($\pm$8.71) 64.72($\pm$6.78) 67.28($\pm$5.96) 64.72($\pm$6.78) 72.09($\pm$12.36) 68.41($\pm$6.75)
NB MFCC 75.63($\pm$7.86) 71.29($\pm$6.93) 71.35($\pm$7.62) 72.67($\pm$8.28) 71.68($\pm$6.37) 72.67($\pm$8.28) 69.91($\pm$11.27) 71.29($\pm$6.93)
MLP MFCC 72.21($\pm$7.85) 60.72($\pm$8.99) 59.08($\pm$8.85) 83.53($\pm$11.27) 68.17($\pm$5.40) 83.53($\pm$11.27) 37.91($\pm$24.21) 60.72($\pm$8.99)
ResNet50 [18] MFCC 75.09($\pm$6.02) 68.74($\pm$6.16) 74.76($\pm$14.74) 64.48($\pm$15.81) 66.42($\pm$8.81) 64.48($\pm$15.81) 73.00($\pm$16.46) 68.74($\pm$6.16)
MobilenetV3 Spec. 78.75($\pm$5.64) 72.59($\pm$5.10) 81.34($\pm$9.68) 61.09($\pm$13.02) 68.39($\pm$7.74) 61.09($\pm$13.02) 84.09($\pm$11.02) 72.59($\pm$5.10)
Dasheng Spec. 85.47($\pm$6.37) 75.18($\pm$8.42) 71.17($\pm$8.58) 88.63($\pm$5.34) 78.49($\pm$5.40) 88.63($\pm$5.34) 61.73($\pm$18.34) 75.18($\pm$8.42)
Table 7: The average classification performance (%) achieved using different enhanced modules and modalities, along with the standard deviation (reported in parentheses), is evaluated across experiments using 10-fold cross-validation on the SYSU-A dataset.
Modality
Quadratic-
-fitting
Diffusion-
-refinement
ROC-AUC Accuracy Precision Recall F-score Sensitivity Specificity
Mean of
Sensi. & Speci.
Audio - - 85.47($\pm$6.37) 75.18($\pm$8.42) 71.17($\pm$8.58) 88.63($\pm$5.34) 78.49($\pm$5.40) 88.63($\pm$5.34) 61.73($\pm$18.34) 75.18($\pm$8.42)
Audio+AGA - - 84.88($\pm$7.66) 75.57($\pm$9.57) 71.73($\pm$9.49) 87.41($\pm$8.14) 78.42($\pm$7.64) 87.41($\pm$8.14) 63.73($\pm$16.92) 75.57($\pm$9.57)
Audio+AGA - 85.82($\pm$4.71) 76.20($\pm$8.86) 73.67($\pm$9.92) 85.03($\pm$6.33) 78.49($\pm$6.57) 85.03($\pm$6.33) 67.37($\pm$17.47) 76.20($\pm$8.86)
Audio+AGA 87.04($\pm$3.69) 78.12($\pm$5.06) 73.32($\pm$6.03) 90.06($\pm$6.52) 80.52($\pm$4.00) 90.06($\pm$6.52) 66.18($\pm$11.35) 78.12($\pm$5.06)
6.6 Evaluation Metrics
To comprehensively evaluate the performance of our proposed system, we employed several widely recognized metrics, including accuracy, precision, sensitivity (recall), specificity, F-score, the mean of sensitivity and specificity, and the area under the receiver operating characteristic curve (ROC-AUC). Together, these metrics, as detailed in Table 7, highlight the robustness and clinical applicability of the proposed system.
7 Experimental Results
7.1 Detection of Vocal Fold Paralysis
To evaluate the effectiveness of our proposed methods, we conduct ablation studies using various system configurations on the clinical dataset. A stratified 10-fold cross-validation approach is employed to ensure robust evaluation. The results are reported as the mean and standard deviation across the 10 folds. The experiments primarily focus on three key aspects:
  1. 1.
    Effectiveness of the pre-trained audio model: Assessing its impact on VFP detection performance comparing with popular SOTA methods.
  2. 2.
    Impact of metric-enhancing modules: Evaluating the contributions of the Quadratic Fitting (QF) and Diffusion Refinement (DR) modules in improving metric extraction, and thus improving the VFP detection performance.
  3. 3.
    Multimodality in VFP detection: Analyzing the benefit of integrating audio and video modalities for enhanced detection performance.
7.1.1 Audio Models vs. VFP Detection Performance
We evaluate the effectiveness of employing pre-trained audio models for VFP prediction by comparing our proposed solution with other widely adopted methods in the field. The results, summarized in Table 7, include RF with MFCC, a model shown to achieve SOTA performance on vocal pathology classification tasks [20]. ResNet50 with MFCC, which has demonstrated superior performance over traditional machine learning models for VFP detection on other datasets [18], is also included. Additionally, we consider commonly used baselines, including SVM, NB, and MLP [20, 18]. For deep learning comparisons, MobileNetV3 [50] is evaluated due to its proven efficiency and effectiveness in image classification tasks.
Our method demonstrates a superior overall balance across these metrics. This is evident from its remarkable F-score and the mean of sensitivity and specificity. Notably, our proposed method achieves an F-score of 78.49%, which outperforms the second-best model by nearly 7%. Furthermore, it maintains a balanced performance in sensitivity and specificity, achieving values of 88.63% and 61.73%, respectively. This results in the highest mean of sensitivity and specificity at 75.18%. Compared to other models, our method achieves high sensitivity while maintaining relatively low specificity, which is crucial for MLVAS in assisting clinicians. This balance reduces the risk of missing unhealthy patients while avoiding the misdiagnosis of an excessive number of healthy patients.
In terms of sensitivity (recall), as shown in Table 7, RF with MFCC achieves an impressive value of 99.27%. However, this is accompanied by a significantly lower precision (53.68%), suggesting a tendency to over-predict positive labels and generate a high number of false alarms. Conversely, MobileNetV3 attains the highest precision of 81.34%, but its sensitivity is relatively low at 61.09%, indicating limited capability in detecting positive cases and potentially overlooking numerous patients genuinely affected by VFP.
Our method leverages fine-tuning of a pre-trained model to achieve an effective balance between precision and recall, thus addressing the limitations observed in other models. This balance underscores the robustness of Dasheng across all performance metrics, positioning it as a reliable and practical approach for VFP detection.
7.1.2 Enhanced Modules vs. VFP Detection Performance
Accurate midline prediction is critical for analyzing left and right vocal fold movements, as it serves as a reference for measuring angle deviations and calculating distances between sample points on the vocal folds and the midline. However, in real-world applications, the laryngeal camera’s position is not always fixed, complicating metric computation. Calibrating the midline by fitting a quadratic function has proven effective for improving midline prediction, leading to better classification performance. The associated results are detailed in Table 7. The system’s native setting serves as a foundational benchmark from which incremental enhancements are realized. The introduction of the proposed QF method improves the specificity from 63.73% to 67.37%, reducing the risk of misdiagnosing healthy patients. Further improvements are achieved by incorporating DR into the glottis segmentation process. As indicated in Table 7, all the metrics show great improvement, except precision and specificity show a slight decrease. ROC-AUC, accuracy, and F-score improve by approximately 2% in absolute terms. Particularly, sensitivity grows from 85.03% to 90.06%, significantly reducing the probability of missing a negative sample. DR is applied only to images where the glottis is not detected by the YOLO-v5 detector (discussed in Section 2.3), as this technique specifically targets false alarms (FAs) in glottis detection (described in Section 6.4). By reducing false alarms, the overall performance of downstream VFP detection improves.
7.1.3 Multimodality vs. VFP Detection Performance
We also investigated the effect of multimodal inputs on VFP detection. When the AGA movement metrics were removed, classification performance decreased on all the metrics, as shown in Table 7. Specifically, specificity fall by an absolute value of 4.45%, and all the other metrics drop by an approximate value of 2% in absolute terms. These results highlight the importance of incorporating multimodality for robust performance.